# Automorphism groups of affine varieties and their Lie algebras

### v1

@@ 9.7. Linearization of reductive group actions
Let $G$ be a reductive linear algebraic group acting effectively on ${\mathbb{A}}^{n}$ with a fixed point $x$. Then the tangent representation of $G_{x}$ on $T_{x}X$ is faithful. Luna’s Étale Slice Theorem guarantees a local linearization of the $G_{x}$-action near $x$. This justifies an expectation that any action of a reductive group on ${\mathbb{A}}^{n}$ can be linearized, that is, is conjugate in $\operatorname{Aut}({\mathbb{A}}^{n})$ to a linear action; see e.g. [Kam79, Conjecture 3.1] and [vdE00, Section 9.4]. However, the latter expectation occurs to be only partially true.
Let us start with reductive subgroups of $\operatorname{Aut}({\mathbb{A}}^{2})$. Every algebraic subgroup of $\operatorname{Aut}({\mathbb{A}}^{2})$ has bounded degree, hence also a bounded length with respect to the free amalgamated product structure, see Theorem 9.3. By a theorem of Serre [Ser03, Ch. I, Sect. 4.3, Theorem 8] (cf. also [KPZ17, Proposition 2.24]) a subgroup of bounded length in an amalgamated free product $A*_{C}B$ is conjugate to a subgroup of one of the factors $A$ and $B$. This leads to the following theorem.
@@
The next corollary was proved by geometric means by T. Igarashi (${}^{\prime}$75, unpublished), see also [Fur83, Theorem 2] and [Miy24, Theorem 2.4.6].
@@
Any action of a finite group on ${\mathbb{A}}^{2}$ is conjugate to a linear action.
In higher dimensions, we start with the actions of commutative reductive groups. The following theorem by Kraft-Schwarz generalizes a similar theorem for algebraic tori by Białynicki-Birula, see [Bia66-67].
Theorem 9.23 ([KrSc92, Theorem VI.3.2(2)]).
Any effective action of a commutative reductive group of dimension $\geq n-1$ on ${\mathbb{A}}^{n}$ is linearizable.
In particular, all maximal tori in $\operatorname{Aut}({\mathbb{A}}^{n})$ are pairwise conjugate and the same holds for tori of dimension $n-1$. The latter fails for affine surfaces and affine threefolds, in general, see Examples 8.13 and 9.20. For the actions of tori of dimension $n-2$ on ${\mathbb{A}}^{n}$ we have the following theorem. Its proof uses a classification of certain smooth affine threefolds diffeomorphic to $\mathbb{R}^{6}$ and non-isomorphic to ${\mathbb{A}}^{3}_{\mathbb{C}}$.
Theorem 9.24 (see [KKMLR97], [KML97], [KoRu97]).
For ${\mathbb{K}}={\mathbb{C}}$, any action of $\operatorname{{\mathbb{G}}_{\mathrm{m}}}$ on ${\mathbb{A}}^{3}_{\mathbb{C}}$ is conjugate to a linear diagonal action.
Theorem 9.25 ([LP21, Corollaire 1.2]).
Any finite subgroup of the tame automorphism group $\mathop{\rm Tame}_{3}$ is linearizable.
Theorem 9.26 ([KaR82]).
Assume that an effective action of the algebraic torus $T=(\operatorname{{\mathbb{G}}_{\mathrm{m}}})^{n-2}$ on ${\mathbb{A}}^{n}$ is ”unmixed,” in the sense that no nontrivial character, together with its inverse, occur as weights of semi-invariant polynomials. Then this action can be linearized.
Theorem 9.27 ([KoRu89]).
Suppose that the algebraic torus $T=(\operatorname{{\mathbb{G}}_{\mathrm{m}}})^{n-2}$ acts effectively on ${\mathbb{A}}^{n}$. If the fixed point set of $T$ has dimension $\geq 1$, then the $T$-action is linearizable.
Next we turn to actions of non-commutative reductive groups. The following theorem is a corollary of Luna’s Étale Slice Theorem; cf. also [BH85, Corollary 10.7].
Theorem 9.28 ([Lun73, Corollare 2 and Remarque on p. 99]).
Any action of a reductive group $G$ on ${\mathbb{A}}^{n}$ with an isolated fixed point and a dense orbit is linearizable.
Theorem 9.29 ([KrPo85], [Pop01] for $n=3$, [Pan84-86] for $n=4$).
For $n\leq 4$, any action of a non-commutative connected semisimple algebraic group on ${\mathbb{A}}^{n}$ is conjugate to a linear action. The same holds for any connected reductive group for $n=3$ and for any such group different from a one- or two-dimensional torus for $n=4$.
Theorem 9.30 ([KrRu14, Theorem A]).
@@
All known examples of non-linearizable actions on ${\mathbb{A}}^{n}$ are obtained from non-trivial algebraic $G$-vector bundles over $G$-modules, following an idea from [KaR82] and [BH85], see also [Kra89], [KrSc92], [MP91] and [MMP91]. Indeed, by the Quillen-Suslin theorem, such a bundle is trivial when ignoring the $G$-action. Hence, its total space is isomorphic to some ${\mathbb{A}}^{n}$. The induced $G$-action on ${\mathbb{A}}^{n}$ is non-linearizable due to the fact that our bundle is non-trivial in the category of $G$-vector bundles.
However, this argument does not work for an abelian reductive group $G$. Indeed, in this case every $G$-vector bundle over a $G$-module is trivial, see [MMP96]. In particular, the following questions remain open, see e.g. [KrSc95] and [Mas03]:
@@
There are several examples of non-linearizable actions of non-abelian finite groups on ${\mathbb{A}}^{n}$. In particular, we have the following theorem.
Theorem 9.33 ([MJ94, Theorem 5.1]).
There is a non-linearizable action of the dihedral group $D_{10}$ on ${\mathbb{A}}^{4}$, and there is a continuous one-parameter family of non-linearizable actions of $D_{14}$ on ${\mathbb{A}}^{4}$.
See e.g. [MP95], [MMP96], [Mas03] and the references therein for further examples of non-linearizable actions of dihedral groups on ${\mathbb{A}}^{n}$ and [FM15] for such examples with an action of the symmetric group $S_{3}$ on ${\mathbb{A}}^{4}$. It seems that no example of this kind is known for a finite group action on ${\mathbb{A}}^{3}$, see e.g. [FM15]. See also [PeRa86] for the Lefschetz fixed-point formula in the context of finite group actions on affine varieties.
The linearization problem is closely related to the Abhyankar–Sathaye problem on linearization of embeddings ${\mathbb{A}}^{k}\hookrightarrow{\mathbb{A}}^{n}$ and on variables in polynomial rings. We send the reader e.g. to [Kal09], [KrRu14], [vdEMV07] and the literature therein for further reading.
@@ 9.8. Solvable and unipotent ind-subgroups: triangulation
The unipotent elements of the de Jonquières group ${\rm Jonq}_{n}$ form a solvable subgroup of derived length $n$. This subgroup is not nilpotent for $n>1$; see [FuKr18, Sect. 15.1 and Lemma 3.2].
@@
We say that a subgroup $G\subseteq\operatorname{Aut}({\mathbb{A}}^{n})$ is triangulable if it is conjugate in $\operatorname{Aut}({\mathbb{A}}^{n})$ to a subgroup of ${\rm Jonq}_{n}$. It is said to be stably triangulable if $G\oplus\operatorname{id}_{{\mathbb{A}}^{k}}$ is triangulable in $\operatorname{Aut}({\mathbb{A}}^{n+k})$ for some $k\geq 0$.
By the Kambayashi-Serre theorem, any unipotent subgroup of $\operatorname{Aut}({\mathbb{A}}^{2})$ is triangulable, see [Ren68] and [Kam79, Theorem 4.3]. This is not any longer true in higher dimensions. However, the next theorem provides a large class of triangulable unipotent subgroups in any dimension.
Theorem 9.36 ([KrZa22, Theorem D]).
@@
Proposition 9.39 ([KrZa22, Proposition 1.4.1]).
@@
An analog of the former assertion fails in ${\mathbb{A}}^{4}$, see [Win90]. For further results in this direction, see e.g. the survey article [Kal09] and [GMM21] for $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-actions on affine varieties, [Fre17] and [vdEKC21] for $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-actions on affine spaces and Hilbert’s 14th problem. See also [Pop17] and the references therein for (stably) birationally triangulable and non-triangulable actions of unipotent groups.
The following question concerns a modified version of stable triangulation.
Question 9.41 ([KrZa22, Question 6]).
Let $G\subseteq\operatorname{Aut}(X)$ be a solvable or nilpotent connected subgroup. Does there exists a closed embedding $X\hookrightarrow{\mathbb{A}}^{n}$ such that $G$ extends to a subgroup of the de Jonquières group $Jonq_{n}$? Is this true if $G$ is algebraic resp., nested ?
The answer is affirmative if $X={\mathbb{A}}^{k}$ and $G\subseteq\operatorname{Aut}({\mathbb{A}}^{k})$ is a unipotent algebraic group, see [KrZa22, Remark 1.4.2(3)]. For instance, the Nagata-Bass $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroup of $\operatorname{Aut}({\mathbb{A}}^{3})$ can be triangulated in ${\mathbb{A}}^{4}$ in the above sense, being non-stably triangulable in the usual sense, cf. Example 9.38.
@@ 10. Tits’ type alternative for automorphism groups
A group $G$ is called virtually solvable resp., virtually nilpotent, etc. if $G$ is a finite extension of a solvable resp. nilpotent, etc. group. The classical Tits alternative says that any linear group $G$, that is, a subgroup of $\operatorname{GL}(n,{\mathbb{K}})$ either is virtually solvable, or contains a free group ${\mathbb{F}}_{2}$ on two generators. For finitely generated subgroups, the Tits alternative holds without any restriction on the base field.
There are several analogs of this theorem for automorphism groups. For instance, the Tits alternative holds for $\operatorname{Aut}({\mathbb{A}}^{2}_{{\mathbb{C}}})$, see [Lam01, Theorem 2.4] or [Lam24, Proposition 20.17].
Notice that Aut($\operatorname{Aut}({\mathbb{A}}^{2}_{{\mathbb{K}}})$ is not isomorphic to a linear group over any transcendental extension of the base field ${\mathbb{K}}$. However, by [Mat23, Theorem A] the codimension 6 normal subgroup
@@
is a linear group over a suitable field extension.
For the proof of the following theorem see [Can11] and [Ure21] for (a), [BFL14, Theorem C] for (b) and [LP22, Main Theorem] for (c).
@@
An analog of the Tits alternative holds for the following groups:
@@
    $\mathop{\rm Tame}(\operatorname{Aut}(\operatorname{SL}(2,{\mathbb{K}})))$ (see Remark 9.8);
@@
    $\mathop{\rm Tame}_{3}=\mathop{\rm Tame}(\operatorname{Aut}({\mathbb{A}}^{3}))$.
@@
    Let $X$ be an affine algebraic surface and $G\subseteq\operatorname{Aut}(X)$ be a subgroup generated by $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroups $U_{1},\ldots,U_{k}$. Then either $G$ contains ${\mathbb{F}}_{2}$ or $G$ is a metabelian unipotent algebraic group.
@@
    Let $X$ be a toric affine variety and $G\subseteq\operatorname{Aut}(X)$ be a subgroup generated by $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroups $U_{1},\ldots,U_{k}$ normalized by the acting torus. Then either $G$ contains ${\mathbb{F}}_{2}$ or $G$ is a unipotent algebraic group.
Similar results also hold for subgroups of $\mathop{\rm Tame}(\operatorname{Aut}(\operatorname{SL}(2,{\mathbb{K}})))$ and $\mathop{\rm Tame}(\operatorname{Aut}({\mathbb{A}}^{3}))$, see [AZ23, Corollary 6.4].

### v2

@@ 9.7. Linearization of reductive group actions
Let $G$ be a reductive linear algebraic group acting effectively on ${\mathbb{A}}^{n}$ with a fixed point $x$. Then the tangent representation of $G_{x}$ on $T_{x}X$ is faithful. Luna’s Étale Slice Theorem [Lun73] guarantees a local linearization of the $G_{x}$-action near $x$. This justifies an expectation that any action of a reductive group on ${\mathbb{A}}^{n}$ can be linearized, that is, is conjugate in $\operatorname{Aut}({\mathbb{A}}^{n})$ to a linear action; see e.g. [Kam79, Conjecture 3.1] and [vdE00, Section 9.4]. However, the latter expectation turns out to be wrong as we will see below in 9.7.
The first results in this context are due to Białynicki-Birula who proved the following in [Bia66-67].
Theorem 9.22.
A faithful action of a torus of dimension $\geq n-1$ on ${\mathbb{A}}^{n}$ is linearizable. In particular, all maximal tori in $\operatorname{Aut}({\mathbb{A}}^{n})$ are pairwise conjugate.
The latter fails for affine surfaces and affine threefolds, in general, see Examples 8.14 and 9.21.
It then took a long time to settle the next open problem, namely the linearizability of a $\mathbb{G}_{\mathrm{m}}$-action on ${\mathbb{A}}^{3}_{{\mathbb{C}}}$. This was done in a series of fundamental papers of Koras and Russell, see e. g. [KoRu97], completed finally in [KKMLR97].
Theorem 9.23.
Any action of $\operatorname{{\mathbb{G}}_{\mathrm{m}}}$ on ${\mathbb{A}}^{3}_{\mathbb{C}}$ is linearizable.
The proof uses a classification of certain Koras-Russell affine threefolds diffeomorphic to $\mathbb{R}^{6}$ and non-isomorphic to ${\mathbb{A}}^{3}_{\mathbb{C}}$, see [KML97].
Certain actions of codimension 2 tori can be linearized also in higher dimensions, see [KamRu82] for (a) and [KoRu89] for (b).
Theorem 9.24.
Assume that the algebraic torus $T=(\operatorname{{\mathbb{G}}_{\mathrm{m}}})^{n-2}$ acts effectively on ${\mathbb{A}}^{n}$. Then this action can be linearized provided one of the following holds:
  1. (a)
    The action is ”unmixed,” that is, no nontrivial character, together with its inverse, occur as weights of semi-invariant polynomials;
  2. (b)
    the fixed point set of $T$ has positive dimension.
Next we look at the case of the affine plane ${\mathbb{A}}^{2}$. Using the amalgamated product structure of $\operatorname{Aut}({\mathbb{A}}^{2})$ and a famous result of Serre’s [Ser03, I.4.3.8] on subgroups of bounded length in an amalgamated product, one gets the following theorem.
@@
In particular, any action of a finite group on ${\mathbb{A}}^{2}$ is conjugate to a linear action. This was first proven by geometric means by T. Igarashi (1975, unpublished), see also [Fur83, Theorem 2] and [Miy24, Theorem 2.4.6].
For actions of reductive groups we have the following consequence of Luna’s Étale Slice Theorem, see [Lun73, III. Corollaire 2]; cf. also [BH85, Corollary 10.7].
Theorem 9.26.
Let a reductive algebraic group $G$ act on an affine variety $X$. Assume that every $G$-invariant function is constant and that $G$ has a smooth fixed point in $X$. Then $X$ is $G$-equivariantly isomorphic to a representation of $G$.
Applied to the actions on ${\mathbb{A}}^{n}$, this gives the following corollary.
Corollary 9.27 ([KrPo85, Proposition 5.1]).
If a reductive group acts on ${\mathbb{A}}^{n}$ with a dense orbit, then the action is linearizable.
For semisimple groups we have the following results, see [KrPo85] and [KrRu14] for $n=3$ and [Pan84-86] for $n=4$.
Theorem 9.28.
For $n\leq 4$, any action of a semisimple algebraic group on ${\mathbb{A}}^{n}$ is linearizable.
In the paper [KrSc92] Kraft and Schwarz study the linearization problem for reductive group actions on smooth acyclic affine varieties $X$ assuming that the quotient $X/\!/G$ has dimension 1. It turns out that there is always a fixed point $x_{0}\in X$ ([KrSc92, Theorem 1]) so that the tangent representation $V:=T_{x_{0}}X$ is the candidate for the linearization. One of the main results is the following ([KrSc92, Theorem 5]).
Theorem 9.29.
$X$ is $G$-isomorphic to $V:=T_{x_{0}}X$ in the following cases:
  1. (1)
    $V$ is a semifree $G$-module;
  2. (2)
    $G$ is a torus;
  3. (3)
    $\dim V^{G^{0}}=1$;
  4. (4)
    $\dim X\leq 3$;
  5. (5)
    $G^{0}$ is semisimple;
  6. (6)
    $V$ is selfdual as a $G^{0}$-module.
As a corollary one gets the next result, cf. [KrRu14, Theorem A].
@@
As for linearizable finite group actions, Kraft and Procesi suggested a conjectural linearization procedure that uses the Reynolds operator. This linearization procedure works indeed in the case of involutions of ${\mathbb{A}}^{n}$ acting with a fixed point, provided the associated Taylor decomposition of the Reynolds average is of bounded length, see [Jur90].
Counterexamples
@@
Many more examples are given in [KrSc92, VII.5. Examples].
All known examples of non-linearizable actions on ${\mathbb{A}}^{n}$ are obtained from non-trivial algebraic $G$-vector bundles over $G$-modules, following an idea from [BH87], see also [Kra89], [KrSc92], [MP91] and [MMP91]. Indeed, by the Quillen-Suslin theorem, such a bundle is trivial when ignoring the $G$-action. Hence, its total space is isomorphic to some ${\mathbb{A}}^{n}$. The induced $G$-action on ${\mathbb{A}}^{n}$ is non-linearizable due to the fact that our bundle is non-trivial in the category of $G$-vector bundles.
However, this method does not work for an abelian reductive group $G$. Indeed, in this case every $G$-vector bundle over a $G$-module is trivial, see [MMP96]. The following questions remain open, see e.g. [KrSc95] and [Mas03]:
@@
There are several examples of non-linearizable actions of non-abelian finite groups on ${\mathbb{A}}^{n}$, for instance, of the dihedral groups $D_{10}$, $D_{14}$ and of of the symmetric group $S_{3}$ on ${\mathbb{A}}^{4}$, see [MJ94, Theorem 5.1] and [FM15]; cf. also [MP95], [MMP96], [Mas03]. It seems that no example of this kind is known for a finite group action on ${\mathbb{A}}^{3}$, see e.g. [FM15]. See also [PeRa86] for the Lefschetz fixed-point formula in the context of finite group actions on affine varieties.
The linearization problem is closely related to the Abhyankar–Sathaye problem on linearization of embeddings ${\mathbb{A}}^{k}\hookrightarrow{\mathbb{A}}^{n}$ and on variables in polynomial rings. We send the reader e.g. to [DG15], [Kal09], [KrRu14], [vdEMV07] and the literature therein for further reading.
@@ 9.8. Solvable and unipotent ind-subgroups: triangulation
The unipotent elements of the de Jonquières group ${\rm Jonq}_{n}$ form a solvable subgroup $\mathcal{J}_{n}$ of derived length $n$, see [FuPo18, Lemma 3.2]. This subgroup is not nilpotent for $n>1$; see Remark 6.3 above.
@@
We say that a subgroup $G\subseteq\operatorname{Aut}({\mathbb{A}}^{n})$ is triangulable if it is conjugate in $\operatorname{Aut}({\mathbb{A}}^{n})$ to a subgroup of ${\rm Jonq}_{n}$. It is said to be stably triangulable if $G\times\operatorname{id}_{{\mathbb{A}}^{k}}$ is triangulable in $\operatorname{Aut}({\mathbb{A}}^{n+k})$ for some $k\geq 0$.
By [Ren68] any unipotent subgroup of $\operatorname{Aut}({\mathbb{A}}^{2})$ is triangulable, cf. Theorem 9.25. This is not any longer true in higher dimensions. However, the next theorem provides a large class of triangulable unipotent subgroups in any dimension.
Theorem 9.35 ([KrZa24, Theorem D]).
@@
Proposition 9.38 ([KrZa24, Proposition 1.4.1]).
@@
An analogue of the former assertion fails in ${\mathbb{A}}^{4}$, see [Win90]. For further results in this direction, see e.g. the survey article [Kal09] and [GMM21] for $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-actions on affine varieties, [Fre17] and [vdEKC21] for $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-actions on affine spaces and Hilbert’s 14th problem. See also [Pop17] and the references therein for (stably) birationally triangulable and non-triangulable actions of unipotent groups.
@@ 10. Tits’ type alternative for automorphism groups
A group $G$ is called virtually solvable resp., virtually nilpotent, etc. if $G$ is a finite extension of a solvable resp. nilpotent, etc. group. The classical Tits alternative says that any linear group $G$, that is a subgroup of $\operatorname{GL}(n,{\mathbb{K}})$, is either virtually solvable, or contains a free group ${\mathbb{F}}_{2}$ on two generators. For finitely generated subgroups, the Tits alternative holds without any restriction on the base field.
There are several similar statements for automorphism groups. For instance, an analogue of the Tits alternative holds for $\operatorname{Aut}({\mathbb{A}}^{2}_{{\mathbb{C}}})$, see [Lam01, Theorem 2.4] or [Lam24, Proposition 20.17].
Notice that Aut($\operatorname{Aut}({\mathbb{A}}^{2}_{{\mathbb{K}}})$ for an infinite field ${\mathbb{K}}$ is not isomorphic to a linear group over any extension of ${\mathbb{K}}$, and the same holds for its de Jonqières subgroup ${\rm Jonq}_{2}$, see [Cor17, Proposition 2.3 and Remark 2.4] and also [Mat23]. However, by [Mat23, Theorem A] the codimension 6 normal subgroup
@@
of Aut($\operatorname{Aut}({\mathbb{A}}^{2}_{{\mathbb{K}}})$ is a linear group over a suitable field extension.
The unipotent triangular subgroup $\mathcal{J}_{n}$ of Aut($\operatorname{Aut}({\mathbb{A}}^{n}_{{\mathbb{K}}})$ over an arbitrary field ${\mathbb{K}}$ is a semidirect product of abelian groups, see [BNS12, Theorem 1]. For $n\geq 3$ this group is nonlinear (it cannot be represented by matrices over any field), see [Sos07, Theorem 2]. The same holds for the tame subgroup $\mathop{\rm Tame}_{n}=\mathop{\rm Tame}(\operatorname{Aut}({\mathbb{A}}^{n}_{{\mathbb{K}}}))$ provided ${\rm char}({\mathbb{K}})=0$ and $n\geq 4$, see [RCS04]. Similar results are also true for the automorphism groups of free associative algebras, see ibid.
For the proof of the following theorem see [Can11] and [Ure21] for (a), [BFL14, Theorem C] for (b) and [LP22, Main Theorem] for (c). In Theorems 10.1 and 10.2 we assume that the base field ${\mathbb{K}}$ is algebraically closed of characteristic zero.
@@
The Tits alternative holds for the following groups:
@@
    $\mathop{\rm Tame}(\operatorname{Aut}(\operatorname{SL}(2,{\mathbb{K}})))$ (see Remark 9.8.3);
@@
    $\mathop{\rm Tame}_{3}=\mathop{\rm Tame}(\operatorname{Aut}({\mathbb{A}}^{3}_{{\mathbb{K}}}))$.
@@
    Let $X$ be an affine algebraic surface over ${\mathbb{K}}$ and $G\subseteq\operatorname{Aut}(X)$ be a subgroup generated by $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroups $U_{1},\ldots,U_{k}$. Then either $G$ contains ${\mathbb{F}}_{2}$ or $G$ is a metabelian unipotent algebraic group.
@@
    Let $X$ be a toric affine variety over ${\mathbb{K}}$ and $G\subseteq\operatorname{Aut}(X)$ be a subgroup generated by Demazure root subgroups $U_{1},\ldots,U_{k}$, see Definition 7.8. Then either $G$ contains ${\mathbb{F}}_{2}$ or $G$ is a unipotent algebraic group.
Similar results also hold for subgroups of $\mathop{\rm Tame}(\operatorname{Aut}(\operatorname{SL}(2,{\mathbb{K}})))$ and $\mathop{\rm Tame}(\operatorname{Aut}({\mathbb{A}}^{3}_{{\mathbb{K}}}))$, see [AZ23, Corollary 6.4].
Acknowledgments. We would like to thank Ivan Arzhantsev, Alexander Perepechko and Andriy Regeta for their comments and references. Our thanks also go to the anonymous referee for useful remarks.